OpenStudy (anonymous):

Find dy/dx and the gradient of the curve at the given value of y: y=(4x-5)ˆ3, y=27

OpenStudy (jack1):

i only know how to do this longhand, hope that's ok?

OpenStudy (anonymous):

Yup it's okay

OpenStudy (jack1):

k, so y=(4x-5)ˆ3 =(4x-5)(4x-5)(4x-5) = 64x^3 - 240x^2 + 300x -125

OpenStudy (anonymous):

Is there any shorter method instead of having to expand it out?

OpenStudy (jack1):

that's the thing, i only know how to do derivatives longhand (expanding first) so dunno, sorry

sam (.sam.):

Just use power and chain rule

OpenStudy (jack1):

y = 64x^3 - 240x^2 + 300x -125 so dy/dx = (64*3)x^2 - (240*2)x + 300

sam (.sam.):

\[y=(4x-5)^3 \\ \\ y'=3(4x-5)^2[\frac{d}{dx}[4x-5]]\]

OpenStudy (campbell_st):

well use the chain rule to differentiate \[y' = 3(4x - 5)^2 \times 4\] so the derivative is \[y' = 12(4x - 5)^2\] to find the gradient at y = 27 you need to find the value of x that makes\[27 = (4x - 5)^3\] then substitute it into the derivative

OpenStudy (anonymous):

When finding x, is the only way to expand it out? Or is there a shorter method?

OpenStudy (campbell_st):

nope take the cube root of both sides so 3 = 4x - 5 solve for x

OpenStudy (anonymous):

@campbell_st Thanks!!